OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at G (1, 6), H (2.5, 4), I (1, 2), K (-0.5, 4). Round your answer to the nearest hundredth when necessary. Select one: a. 5 units b. 12.5 units c. 7.5 units d. 10 units

OpenStudy (anonymous):

@eliassaab @abb0t

OpenStudy (anonymous):

its between c and d I just cant deside

OpenStudy (anonymous):

Use the distance formula to find the distance of each side of the quadrilateral and then add the lengths of each side to find the perimeter.

OpenStudy (anonymous):

like this The length of GH is [2.5 - (1)] = 1.5 The length of HI is √[(1-2.5)^2 + (2-4)^2] = √[2.25+4] = 2.5 The length of IK is [1 - (-0.5)] = 1.5 The length of GK is √[[1 - (-0.5)]^2 + (6-4)^2] = √[1.22+4] = 2.29 The perimeter is the sum: 1.5+2.5+1.5+2=7.5 units

OpenStudy (anonymous):

or like this GH=2.5 IK=2.5 IH=2.5 GK=2.5 Perimeter =10

OpenStudy (anonymous):

GH=2.5 IK=2.5 IH=2.5 GK=2.5 perimeter = 10